find an equation of the form f(x)=mx+b:
f(0)=5 and f(2)=7

Answer by london maths tutor(243)

f(x) = mx+b
f(0) = m(0) + b = b = 5
Therefore b = 5
f(2) = m(2)+ 5 = 7
m(2) = 7 - 5
m(2) = 2
m = 1
Therefore equation : f(x) = x + 5